HUBBLE SPACE TELESCOPE DAILY REPORT #1536 PERIOD COVERED: 0000Z (UTC) 12/13/95 - 0000Z (UTC) 12/14/95 Daily Status Report as of 348/0000Z 1.0 ACTIVITIES SCHEDULED AND ACCOMPLISHED: 1.1 Completed WF/PC-2 5897 (Ages For the Oldest Star Clusters in the = Large Magellanic Cloud) The WF/PC-2 was used to make observations of star cluster NGC-2210 = which is known to be one of the oldest clusters in the Large Magellanic = Cloud. From these data, a color magnitude diagram will be constructed = that extends below the main sequence turnoff. The observations were = completed as planned, and no anomalies were noted. 1.2 Completed WF/PC-2 5999 (The Nuclear Structure of S0 Galaxies: A PC = Imaging Survey) The WF/PC-2 was used to make observations of S0 class galaxy = NGC-3599 in order to study its nuclear structure. This is a part of a = large survey of S0, E/S0, and S0/a class galaxies. The targets will cover = a wide range in luminosity and environment. The observations were = completed as planned, and no anomalies were noted. 1.3 Completed WF/PC-2 5996 (The Morphology of Field Galaxies at High = Redshifts) The WF/PC-2 was used to make observations of several field = galaxies (Field-0302) in order to study their morphologies. The = observations were completed as planned, and no anomalies were reported. 1.4 Completed WF/PC-2 5998 (The Evolution of Massive Stars: Closing the = Loop Observationally) The WF/PC-2 was used to make observations of (OB Class) massive = star M-31-OB69 (field 6) which is located in the Andromeda galaxy. The = goal is to study the evolution of this class of massive stars. The = observations were executed as scheduled, and no anomalies were noted. 1.5 Completed WF/PC-2 5972 (Determination of the Extragalactic Distance = Scale) The WF/PC-2 was used to make observations of the galaxy NGC-3198 = looking for standard candle Cepheid variable stars. The purpose of this = observations is to help estimate the Hubble constant with an accuracy of = 10%. The observations were executed as scheduled, and no anomalies were = noted. 1.6 Completed WF/PC-2 6102 (Chemical Evolution and History of Star = Formation in the Large Magallanic Cloud) The WF/PC-2 was used to make observations of the "Bar" region = (NGC-2682) in the Large Magellanic cloud to study the chemical evolution = and the history of star formation in that region. The goal is to observe = individual field stars near the main sequence turnoff in order to study = the large scale process that control star formation and nucleosynthesis in = galaxies. The observations were executed as planned, and no anomalies = were reported. 1.7 Completed WF/PC-2 5845 (Polarization of Disks around T-Tauri) The PC and polarization filter POLQ were used to image the = nebulosity surrounding the T Tauri star HL-Tau. These data will test = various models for scattered light from central sources. Systematic = errors will be avoided by using one polarizer and allowing the spacescaft = roll to provide the rotation by observing the target three times at 120 = day intervals. The observations were executed as scheduled, and no = problems were reported. 1.8 Completed GHRS/1 5945 (Diffuse Helium Absorption as a Probe of = Quasar History) The High Resolution Spectrograph (Side One) was used to make = observations of diffuse helium absorption lines in the spectra of quasar = Q0302-0019 to confirm the detection of intergalactic helium by the FOC. = The observations were completed as planned and no anomalies were reported. 2.0 FLIGHT OPERATIONS SUMMARY: 2.1 Guide Star Acquisitions: Scheduled Acquisitions : 6 Successful : 6 Scheduled Reacquisitions : 9 Successful : 9 2.2 FHST Updates: Scheduled : 13 Successful : 13 2.3 Operations Notes: At approximately 347/2010Z, the operations team began to observe = symptoms of TAC problems. Eventually, a switch was made from the primary = rail to the backup rail at 347/2153Z. As a result, operations was unable = to utilize a planned Madrid DSN pass. During this time frame the VDS = communications system went down for about 30 minutes. MOR operations moved to the SAMS backup facility for the scheduled = weekly checkout. The DOC was unable to switch over the command lines = until 348/0005Z due to an unknown problem as the NCC. Operations moved = back to the MOR without further incident as 348/0345Z. There was no impact to the on-going science programs as a result of = any of these problems. 3.0 SIGNIFICANT FORTHCOMING EVENTS: Nothing at this time. /DMH